OpenStudy (anonymous):

INTEGRAL OF XSIN^2XDX

OpenStudy (turingtest):

Use integration by parts, and\[\sin^2x=\frac12-\frac12\sin(2x)\]

OpenStudy (anonymous):

U dont use the sin\[\sin ^{2}x=1/2(1+\cos 2x)?\]

OpenStudy (turingtest):

sorry\[\sin^2x=\frac12(1-\cos(2x)\]